The following information was obtained from the unadjusted trial balance of the Pie Corporation for the year ending December 31, 2020:

Accounts Receivable P 2,500,000dr

Allowance for Doubtful Accounts 34,000dr

Sales (90% are on credit) 8,400,000cr

Instructions: Determine the following:

1. Doubtful account expense assuming that 6% of the accounts receivable are estimated to be uncollectible.

2. Amortized cost of accounts receivable.
Answer #1

Answer:

Requirement 1

Doubtful account expense = Accounts Receivable x 6% = 2500000 x 6% = 150000 + 34000 = 184000

Allowance for Doubtful Accounts = 34000 Dr.

Doubtful account expense Dr. 184000 ; Allowance for Doubtful Accounts Cr. 184000

Requirement 2

Amortized cost of accounts receivable.= Accounts Receivable - Allowance for Doubtful Accounts = 2500000 - 184000 = P 2316000